OpenStudy (bj11684):

A marina is in the shape of a coordinate grid. Boat A is docked at (2.4, −3) and Boat B is docked at (−3.4, −3). The boats are ____ units apart. A. 5.8 B. 6.4 C. 8.8 D. 9.4

OpenStudy (faiqraees):

d=√(x1-x2)²+(y1-y2)²

OpenStudy (bj11684):

I do not understand?

OpenStudy (welshfella):

the 2 points can be written as (x1,y1) and (x2 , y2) so in your case you can write x1 = 2.4 , x2 = -3.4 , y1 = -3 and y2 = -3 plug these into the above formula and work it out so we have sqrt [( 2.4 - (-3.4)^2 + (-3-(-3)^2]

OpenStudy (welshfella):

= sqrt (33.64 - 0) = sqrt 33.64

OpenStudy (welshfella):

use your calculator to find square root of 33.64 and you have your answer

OpenStudy (welshfella):

that formula give you d, the distance between any 2 points on the coordinates plane

OpenStudy (welshfella):

actually I just realised we could have done this one much more quickly by subtracting the 2 x coordinates because the y coordinates are both - 3 so d = 2.4 - (-3.4) = 2.4 + 3.4 = what?

OpenStudy (welshfella):

OK?